OpenStudy (anonymous):

what is the sum of the interior angles of a 16- sided polygon

OpenStudy (anonymous):

For an N-sided polygon, the sum of the interior angles is: N*(180° - 360°/N) or just N*180° - 360°. See related link. So for N=16, the sum is 2520

OpenStudy (anonymous):

2520

OpenStudy (anonymous):

so it would be 16(180 - 360) :/

OpenStudy (anonymous):

no jodas. lol jk

OpenStudy (anonymous):

The sum of ALL the measures will equal 2520

OpenStudy (anonymous):

yeah but is that how i set it up ???

OpenStudy (anonymous):

S = 180(n-2) S = 180(16-2) S = 180(14)

OpenStudy (anonymous):

ohhh i c thanks :D

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

yez, yuo